OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (-7,-1) and (-3,-8) The equation of the line in slope-intercept form is y=______ How do I do this?

OpenStudy (callisto):

slope = [-1-(-8)]/[-7-(-3)] = -7/4 equation is y-(-1) = -7/4[x-(-7)] y= -7/4x - 53/4

OpenStudy (anonymous):

Thank you, that helped. But how did you get the 53?